OpenStudy (anonymous):

I need help solving this proof. (It'll be in my next reply so I can do the formatting on it)

OpenStudy (anonymous):

\[4|3^{2n-1} +1\] for integer >= 1

OpenStudy (anonymous):

There are two ways you can prove this: (1) using congruences and modular arithmetic, or (2) mathematical induction. Are you familiar with either of these methods?

OpenStudy (anonymous):

I am familiar with induction. I have used it to solve similar proofs but they are usually much less complex stuff like basic summation formulas.

OpenStudy (anonymous):

And here is what I have so far. 3^(2k-1) + 1 = 4m 3^(2k+1-1) + 1 = 4m

OpenStudy (anonymous):

The first thing you want to do is prove the base case \(n=1\). Can you do that? Then you want to assume for some \(k>1\) that \(\large 4\mid 3^{2k-1}+1\implies 3^{2k-1}+1=4m\). Then your objective is to show that you can rewrite \(\large 3^{2(k+1)-1}+1 = 3^{2k+1}+1\) as a multiple of \(4\). So we note that \[\large \begin{aligned}3^{2k+1}+1 &= 3^2\cdot 3^{2k-1}+1\\ & = 9\cdot 3^{2k-1}+1\\ &= (8+1)\cdot 3^{2k-1}+1\\ &= \ldots\end{aligned}\] Can you see how to proceed from here? :-)

OpenStudy (anonymous):

Yeah base case seems simple enough you eventually get 4|4. Do you then take (8+1)? Leaving you with 4(2+1)? Is that it? Is that the whole thing?

OpenStudy (anonymous):

That should be one fourth and not 1.

OpenStudy (anonymous):

Well I guess since 4 * anything makes that something divisible by 4 that proves that the right side is divisible by 4. But why did we start with the original equation and not the k+1?

OpenStudy (anonymous):

Though I suppose you could do... (8+1)*(3^2k) 4(2+1/4)*(3^2k) = 4m Is that right?

OpenStudy (anonymous):

And the m on the right side would be.....(2+1/4)*(3^2k)...

OpenStudy (anonymous):

Though that can't be right because 2 + 1/4 is not an integer so it's not an integer by closure. So yeah I don't know what to do.

OpenStudy (anonymous):

continuing where I left off, we see that \[\large \begin{aligned}(8+1)\cdot 3^{2k-1}+1 &= 8\cdot 3^{2k-1} + \underbrace{3^{2k-1}+1}_{=4m\text{ by ind. hypothesis}}\\ &= 8\cdot 3^{2k-1} + 4m\\ &= 4\underbrace{(2\cdot 3^{2-1}+m)}_{\in\Bbb{z};\text{ call it $a$}}\\ &= 4a\end{aligned}\] Thus, \(\large 4\mid 3^{2k+1}+1\) and the inductive step is complete. Therefore \(\large 4\mid 3^{2n-1}+1\). Does this make sense? :-)

OpenStudy (anonymous):

The last under brace is supposed to read "\(\in\Bbb{Z}; \text{ call it $a$}\)"

OpenStudy (anonymous):

How did you go from 8+1 to 8 *? I am assuming you multiplied 1 with what was on the right side and then added it?

OpenStudy (anonymous):

Well \(\large (8+1)\cdot 3^{2k-1} = 8\cdot 3^{2k-1} + 3^{2k-1}\) by the distributive property.

OpenStudy (anonymous):

And then you set 3^2k-1 + 1 to 4m? Can you explain why you are able to do that?

OpenStudy (anonymous):

I noticed that I had two typos in that equation block; here's the corrected form. \[\large \begin{aligned}(8+1)\cdot 3^{2k-1}+1 &= 8\cdot 3^{2k-1} + \underbrace{3^{2k-1}+1}_{=4m\text{ by ind. hypothesis}}\\ &= 8\cdot 3^{2k-1} + 4m\\ &= 4\underbrace{(2\cdot 3^{2k-1}+m)}_{\in\Bbb{Z};\text{ call it $a$}}\\ &= 4a\end{aligned}\]

OpenStudy (anonymous):

We made the assumption that for some \(\large k\), we have \(\large \color{red}{3^{2k-1}+1=4m}\); this is known as the inductive hypothesis. Whenever you try to show the statement is true for \(\large k+1\), you need to use the inductive hypothesis at some point. So when you rewrite \[\large 9\cdot 3^{2k-1}+1 = (8+1)\cdot 3^{2k-1}+1 = 8\cdot 3^{2k-1}+\color{red}{3^{2k-1}+1} \] we see the term that appears in our inductive hypothesis pop up, so that's where you make the substitution. Does this clarify things?

OpenStudy (anonymous):

Okay I understand why you are able to place that into 4m. Why did we assume 3^2k-1 = 4m and not 3^2k+1-1 = 4m

OpenStudy (anonymous):

I forgot a plus 1 in that last equation.

OpenStudy (anonymous):

First note that \(\large\displaystyle 4\mid 3^{2n-1}+1 \implies 3^{2n-1}+1=4m\). So then, if you let \(\large n=k\), you have \(\large 3^{2k-1}+1=4m\) and the objective of the induction is to show that \(\large 3^{2(k+1)-1}+1=4a\). If you want to start with \(\large 3^{2k+1}+1=4m\), then the objective of induction is to show that \(\large 3^{2(k+2)-1}+1 =4a \implies 3^{2k+3}+1=4a\). In general, when you prove something with mathematical induction, we assume that it's true for k, and then show it's true for k+1. You could alternately assume it's true for k+1, but then you'd have to show it's true for k+2. Does this make sense?

OpenStudy (anonymous):

Ah I gotcha. I had just seen other problems done by doing some stuff with whatever your term was plus 1.